Summary of Apr. 9 to Apr. 16th

• Had 148 scheduled hours. Got 101 ABU hours (68%) and 120 hours of experiment ready (82%)

– Major problems:

∗ Target lost turbopump on scattering chamber ∗ Lost LHe and LN in dipole, LHe in Q3 and

quench Moller solenoid.

∗ Various quad trips

– Accomplishments ∗ Took more LD2 and LH2 data at Q2= 1.47 ∗ Took more LH2 data at Q2=0.45

• Plans for this week.

– Today change to 1.134/pass and run Hall C at 2 pass. Expect 83% of the injector polarization in Hall.

– Tuesday Day shift do Moller tune up.

– Friday beam stops at noon for Open House. Beam resumes Sunday 8am.

– Run until April 26th.